# via Dodecaneso 35, 16146, Genova, Italy

### v1

@@ 2.2 Towards a random CCZ construction
Let $F$ and $G$ be two CCZ-equivalent functions as in Definition 2.1. We can write the affine bijection $\mathcal{A}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ as
$\mathcal{A}(x,y)=\mathcal{L}(x,y)+(a,b),$
with $a\in{\mathbb{F}}_{q}^{n}$ and $b\in{\mathbb{F}}_{q}^{m}$ and $\mathcal{L}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ linear bijection. Thus, $\mathcal{L}$ maps the graph of $F$ to the graph of $G^{\prime}$, with $G^{\prime}(x)=G(x+a)+b$. Hence, up to a translation of the input and the output, we can consider directly the linear bijection $\mathcal{L}$. Then, we can write $\mathcal{L}$ as a matrix composed by four linear maps,
$\mathcal{L}(x,y)=\begin{bmatrix}A_{1}&A_{2}\\ A_{3}&A_{4}\end{bmatrix}\cdot\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}A_{1}(x)+A_{2}(y)\\ A_{3}(x)+A_{4}(y)\end{bmatrix}=\begin{bmatrix}L_{1}(x,y)\\ L_{2}(x,y)\end{bmatrix}.$
Recall that $\mathcal{L}(x,F(x))=(x^{\prime},G(x^{\prime}))$. Set
$F_{1}(x)=L_{1}(x,F(x))=A_{1}(x)+A_{2}(F(x))$
@@
    A random linear map $L_{1}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}$ such that $L_{1}(x,F(x))$ is a bijection;
@@
Clearly, $L_{1}$ (and consequentially $L_{2}$) strongly depends on the choice of the initial function $F$. So, differently from the affine and the EA equivalence, it appears to be not easy to provide a general construction method for a random CCZ-equivalent function.
@@ 3 The twisting
To study a possible way to construct a random CCZ-equivalent map, we consider a particular instance of CCZ equivalence, introduced for the case $q=2$ by Canteaut and Perrin with the name of twisting. Indeed, in [CP19, Theorem 3] the authors showed that any two CCZ-equivalent functions are connected via the following 3 steps: EA transformation, $t$-twisting, EA transformation.
@@ 3.1 Definition of twisting
$M_{t}(\mathcal{G}_{F})=$ $M_{t}\cdot\begin{bmatrix}x\\ y\\ T(x,y)\\ U(x,y)\end{bmatrix}=\begin{bmatrix}T(x,y)\\ y\\ x\\ U(x,y)\end{bmatrix}.$
To clarify the notation, here and in the rest of the paper we identify $x$ with the $t$ variables $x_{1},\dots,x_{t}$, and $y$ with the $n-t$ variables $y_{1},\dots,y_{n-t}$. The following observation is important.
Remark 3.2.
We have that $M_{t}(\mathcal{G}_{F})$ is the graph of some function (we equivalently say that $M_{t}$ is admissible for $F$) if $T(x,y)=T_{y}(x)$ is a bijection for every fixed $y\in{\mathbb{F}}_{q}^{n-t}$. Then, the CCZ-equivalent function $G$ ($\mathcal{G}_{G}=M_{t}(\mathcal{G}_{F})$) has the form $G(x,y)=(T_{y}^{-1}(x),U_{T_{y}^{-1}(x)}(y))$. Notice that, if the degree of $T_{y}^{-1}(x)$ is $d$ (i.e. the algebraic degree with respect to both $x$ and $y)$, then the degree of $U(T_{y}^{-1}(x),y)$ is at most $2d$.
@@ 3.2 Twisting of quadratic functions
As mentioned in the introduction, the majority of multivariate cryptographic schemes appearing in the literature deals with quadratic functions. Therefore, we consider the case when the central map is a quadratic function and we present a preliminary study of the twisting for quadratic maps.
Let $F$ be a quadratic function admitting a $t$-twist. We keep the notation introduced in §3.1 and we write $F(x,y)=(T(x,y),U(x,y))$. Since $F$ is quadratic, both functions $T$ and $U$ have degree at most 2. Moreover, by Remark 3.2 we have that for every $y\in{\mathbb{F}}_{q}^{n-t}$, $T(x,y)=T_{y}(x)$ is invertible. Then, the bijection $T_{y}(x)$ is either affine or quadratic. We deal with both cases separately.
@@ 3.2.1 $T$ affine
We assume that $T$ is affine, that is $T(x,y)=\ell(x)+\phi(y)$ with $\ell:{\mathbb{F}}_{q}^{t}\rightarrow{\mathbb{F}}_{q}^{t}$ linear bijection and $\phi:{\mathbb{F}}_{q}^{n-t}\rightarrow{\mathbb{F}}_{q}^{t}$ affine transformation. In this case, the map is easily invertible as $T_{y}^{-1}(x)=\ell^{-1}(x)-\ell^{-1}(\phi(y))$. However this map will produce a function $G$ that is EA-equivalent to $F$. Indeed, $G=F\circ A$, with $A$ the affine bijection of the form $A(x,y)=(\ell^{-1}(x)-\ell^{-1}(\phi(y)),y)$. So, we are not interested in this case.
@@ 3.2.2 $T$ quadratic
The general case $T_{y}(x)$ quadratic is quite difficult to analyse. By Remark 3.2, to apply a CCZ transformation, we need the map $T(x,y)=T_{y}(x)$ to be invertible for every fixed $y$. A possible way to achieve this is to consider $T(x,y)$ as a function of the form
@@
The map $T_{y}^{-1}(x)$ has degree at most two. Thus, Remark 3.2 implies that the map $G$ constructed with the $t$-twist has degree at most four.
@@ 4 Multivariate CCZ Scheme
We propose a scheme where the central map is hidden by an application of a CCZ transformation. We present two versions of this proposal at increasing level of details. First, we present a generic instance of the scheme which can be applied to any central map admitting a $t$-twist (§4.1). Then, we present some restrictions on the choice of the quadratic secret map used (§4.2.1 and §4.2.2), summarized in a concrete instance with a specific choice for the central map (§4.2.3).
@@ 4.1 Generic CCZ Scheme
Finally, we construct the public map as $G_{pub}=A_{1}\circ G\circ A_{2},$ for $A_{1},A_{2}$ random affine bijections of ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively. In the secret key, we need to store the information needed to invert the map $G_{pub}$, consisting of $\langle A_{1},A_{2},t,T,U\rangle$. Equivalently, we can directly consider $\langle A_{1}^{-1},A_{2}^{-1},t,T,U\rangle$.
@@ 4.1.1 Proposal as an encryption scheme
Encryption. The sender encrypts the message $m\in{\mathbb{F}}_{q}^{n}$ by evaluating the public key $G_{pub}$ and sends $c=G_{pub}(m)\in{\mathbb{F}}_{q}^{m}$.
Decryption. From $c\in{\mathbb{F}}_{q}^{m}$, the receiver has to compute its preimages, i.e. the set of solutions $\bar{m}$ such that $G_{pub}(\bar{m})=c$. This can be obtained in the following way.
@@
    Given $c$, compute $c^{\prime}=A_{1}^{-1}(c)$. To simplify the notation, for $c=G_{pub}(m)$, call $m^{\prime}=A_{2}(m)$, so $c^{\prime}=G(m^{\prime})$. Moreover, write $c^{\prime}$ as $(c^{\prime}_{T},c^{\prime}_{U})\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{m-t}$ and similarly consider $m^{\prime}=(x,y)\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{n-t}$. So it holds
@@
    For every possible pair of solutions $(\bar{x},\bar{y})$, compute $\bar{m}=A_{2}^{-1}(\bar{x},\bar{y})$.
@@ 4.1.2 Proposal as a signature scheme
In a signature scheme, a sender produces a signature for a document, knowing the secret information. The receiver checks the validity of the signature for the received document.
@@
    We need to find $x\in{\mathbb{F}}_{q}^{t}$ and $y\in{\mathbb{F}}_{q}^{n-t}$ such that $G(x,y)=w$. So $T_{y}^{-1}(x)=w_{T}$ and $U(T_{y}^{-1}(x),y)=w_{U}$.
@@
Verification. The receiver wants to verity that the signature $(x^{\prime},y^{\prime})$ is valid for the document $d$.
@@
The correctness for the encryption scheme $G_{pub}(\bar{m})=c$ is verified in the same way.
@@
Regarding the function $U(x,y)$, different conditions must be satisfied. If we want to use $F$ as a central map for a signature scheme, we need that for any possible $(a,b)\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{m-t}$, the system $U(a,y)=b$ has always a solution. Instead, for the encryption scheme, we need the system to have few solutions. Indeed, in this case, the receiver has to consider all possible pre-images of the map $G_{pub}$ and find the correct one.
@@ 4.2 Our proposal
Assume that $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$ is a quadratic function which admits a $t$-twist, so $F(x,y)=(T(x,y),U(x,y))$ where both $T$ and $U$ have at most degree 2. In what follows we present our proposal for the choice of the maps $T$ an $U$.
@@ 4.2.1 The choice of $T$
leading to a $t$-twist of the form $G(x,y)=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. Observe that with this choice of $T$, to compute $\bar{x}=T_{\bar{y}}(c_{T})$ simply corresponds to computing $\bar{x}=c_{T}+{\mathfrak{q}}(\bar{y})$.
@@ 4.2.2 The choice of $U$
with $\{z_{j}:j\in V\}=\{x_{1},\ldots,x_{t},y_{1},\ldots,y_{s}\}$ and $\{z_{j}:j\in O\}=\{y_{s+1},\ldots,y_{n-t}\}$ respectively the sets of vinegar and oil variables, and with the coefficients $\alpha_{jk}^{(i)},\beta_{jk}^{(i)},\gamma_{j}^{(i)},\delta^{(i)}$ randomly chosen over ${\mathbb{F}}_{q}$. Notice that, fixed the vinegar variables, the system is linear in the oil variables, hence it is easy to solve, for example with a simple Gaussian reduction. The legitimate user can get the preimages of $U$ with respect to $y$ (fixed $x$) using classical techniques from OV systems [DY09, DPS20]. Notice that the easiest system is obtained for $s=0$, however from the analysis presented in §6, this is not a good choice.
@@ 4.2.3 Pesto
Definition 4.2 (Pesto primitive).
@@
Set $G=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. Then the map $G_{pub}=A_{1}\circ G\circ A_{2}$ is the public key, and $\langle A_{1},A_{2},{\mathfrak{q}},U\rangle$ (and eventually $t$) constitutes the secret key.
We do not suggest specific values for the size of the parameters. However, in the following remark we stress the role of $s$.
@@
Hence, when choosing the parameters $n,m,t,s$ one has also to consider if the linear system to solve, $m-t$ equations in $n-t-s$ variables, should be:
@@
Finally, based on the security analysis performed in §6, we also recommend that $t$ is around $n/3$.
@@ 4.2.4 Toy Example
therefore $T(x,y)=x+{\mathfrak{q}}(y)$. The map $U$ is an OV system of 2 equations with $x_{1},x_{2},y_{1}$ oil variables and $y_{2},y_{3}$ vinegar variables. Hence we consider

### v2

@@ 2.2 Towards a random CCZ construction
Let $F$ and $G$ be two CCZ equivalent functions as in Definition 2.1. We can write the affine bijection $\mathcal{A}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ as
$\mathcal{A}(z,w)=\mathcal{L}(z,w)+(a,b),$
with $a\in{\mathbb{F}}_{q}^{n}$ and $b\in{\mathbb{F}}_{q}^{m}$ and $\mathcal{L}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}$ linear bijection. Thus, $\mathcal{L}$ maps the graph of $F$ to the graph of $G^{\prime}$, with $G^{\prime}(z)=G(z+a)+b$. Hence, up to a translation of the input and the output, we can consider directly the linear bijection $\mathcal{L}$. Then, we can write $\mathcal{L}$ as a matrix composed by four linear maps,
$\mathcal{L}(z,w)=\begin{bmatrix}A_{1}&A_{2}\\ A_{3}&A_{4}\end{bmatrix}\cdot\begin{bmatrix}z\\ w\end{bmatrix}=\begin{bmatrix}A_{1}(z)+A_{2}(w)\\ A_{3}(z)+A_{4}(w)\end{bmatrix}=\begin{bmatrix}L_{1}(z,w)\\ L_{2}(z,w)\end{bmatrix}.$
Recall that $\mathcal{L}(z,F(z))=(z^{\prime},G(z^{\prime}))$. Set
$F_{1}(z)=L_{1}(z,F(z))=A_{1}(z)+A_{2}(F(z))$
@@
    A random linear map $L_{1}:{\mathbb{F}}_{q}^{n}\times{\mathbb{F}}_{q}^{m}\rightarrow{\mathbb{F}}_{q}^{n}$ such that $L_{1}(z,F(z))$ is a bijection;
@@
Clearly, $L_{1}$ (and consequentially $L_{2}$) strongly depends on the choice of the initial function $F$. So, differently from the affine and the EA equivalence, it appears to be not easy to provide a general construction method for a random CCZ equivalent function.
@@ 3 The twisting
To study a possible way to construct a random CCZ equivalent map, we consider a particular instance of CCZ equivalence, introduced for the case $q=2$ by Canteaut and Perrin with the name of twisting. Indeed, in [CP19, Theorem 3] the authors showed that any two CCZ equivalent functions are connected via the following 3 steps: EA transformation, $t$-twisting, EA transformation.
@@ 3.1 Definition of twisting
$M_{t}((x,y),F(x,y))=$ $M_{t}\cdot\begin{bmatrix}x\\ y\\ T(x,y)\\ U(x,y)\end{bmatrix}=\begin{bmatrix}T(x,y)\\ y\\ x\\ U(x,y)\end{bmatrix}.$ (2)
To clarify the notation, here and in the rest of the paper we identify $x$ with the $t$ variables $x_{1},\dots,x_{t}$, and $y$ with the $n-t$ variables $y_{1},\dots,y_{n-t}$.
From (2) we obtain immediately the following important result.
Theorem 3.2.
Let $F$ be as in (1) and $M_{t}$ be as in Definition 3.1. If $T_{y}:{\mathbb{F}}_{q}^{t}\to{\mathbb{F}}_{q}^{t}$ is a bijection for any $y\in{\mathbb{F}}_{q}^{n-t}$, then $M_{t}(\mathcal{G}_{F})$ is a graph of a function $G$ (i.e. $\mathcal{G}_{G}=M_{t}(\mathcal{G}_{F})$) given by $G(x,y)=(T_{y}^{-1}(x),U(T_{y}^{-1}(x),y))$. Moreover, if the degree of $T_{y}^{-1}(x)$ is $d$ (i.e. the algebraic degree with respect to both $x$ and $y)$ and the algebraic degree of $U$ is $d^{\prime}$, then the degree of $U(T_{y}^{-1}(x),y)$ is at most $d\cdot d^{\prime}$.
@@ 3.2 Twisting of quadratic functions
As mentioned in the introduction, the majority of multivariate cryptographic schemes appearing in the literature deal with quadratic functions. Therefore, we consider the case when the central map is a quadratic function and we present a preliminary study of the twisting for quadratic maps.
Let $F$ be a quadratic function admitting a $t$-twist. We keep the notation introduced in §3.1 and we write $F(x,y)=(T(x,y),U(x,y))$. Since $F$ is quadratic, both functions $T$ and $U$ have degree at most 2. By Theorem 3.2, we require that for every $y\in{\mathbb{F}}_{q}^{n-t}$, $T(x,y)=T_{y}(x)$ is invertible. Then, the bijection $T_{y}(x)$ is either affine or quadratic. We deal with both cases separately.
@@ 3.2.1 $T$ affine
We assume that $T$ is affine, that is $T(x,y)=\ell(x)+\phi(y)$ with $\ell:{\mathbb{F}}_{q}^{t}\rightarrow{\mathbb{F}}_{q}^{t}$ linear bijection and $\phi:{\mathbb{F}}_{q}^{n-t}\rightarrow{\mathbb{F}}_{q}^{t}$ affine transformation. In this case, the map is easily invertible as $T_{y}^{-1}(x)=\ell^{-1}(x)-\ell^{-1}(\phi(y))$. However, this map will produce a function $G$ that is EA equivalent to $F$. Indeed, $G=F\circ A$, with $A$ the affine bijection of the form $A(x,y)=(\ell^{-1}(x)-\ell^{-1}(\phi(y)),y)$. So, we are not interested in this case.
@@ 3.2.2 $T$ quadratic
The general case $T_{y}(x)$ quadratic is quite difficult to analyse. By Theorem 3.2, to apply a CCZ transformation, we need the map $T(x,y)=T_{y}(x)$ to be invertible for every fixed $y$. A possible way to achieve this is to consider $T(x,y)$ as a function of the form
@@
With $T$ as in (3), the map $T_{y}^{-1}(x)$ has degree at most two. Thus, Theorem 3.2 implies that the map $G$ constructed with the $t$-twist has degree at most four.
@@ 4 Multivariate CCZ Scheme
We propose a scheme where the central map is hidden by an application of a CCZ transformation. We present two versions of this proposal at increasing level of details. First, we present a generic instance of the scheme which can be applied to any central map admitting a $t$-twist (§4.1). Then, we present some restrictions on the choice of the quadratic secret map used (§4.2.1 and §4.2.2), summarized in a concrete instance with a specific choice for the central map (§4.2.3). Finally, after conducting a preliminary security analysis in the following sections, we will present concrete parameters for this proposal in §6.7.
@@ 4.1 Generic CCZ Scheme
Finally, we construct the public map as $G_{pub}=A_{1}\circ G\circ A_{2},$ for $A_{1},A_{2}$ random affine bijections of ${\mathbb{F}}_{q}^{m}$ and ${\mathbb{F}}_{q}^{n}$ respectively. In the secret key, we need to store the information needed to invert the map $G_{pub}$, consisting of $\langle A_{1},A_{2},T,U\rangle$. Equivalently, we can directly consider $\langle A_{1}^{-1},A_{2}^{-1},T,U\rangle$.
@@ 4.1.1 Proposal as an encryption scheme
Encryption. The sender encrypts the message $z\in{\mathbb{F}}_{q}^{n}$ by evaluating the public key $G_{pub}$ and sends $c=G_{pub}(z)\in{\mathbb{F}}_{q}^{m}$.
Decryption. From $c\in{\mathbb{F}}_{q}^{m}$, the receiver has to compute its preimages, i.e. the set of solutions $\bar{z}$ such that $G_{pub}(\bar{z})=c$. This can be obtained in the following way.
@@
    Given $c$, compute $c^{\prime}=A_{1}^{-1}(c)$. To simplify the notation, for $c=G_{pub}(z)$, call $z^{\prime}=A_{2}(z)$, so $c^{\prime}=G(z^{\prime})$. Moreover, write $c^{\prime}$ as $(c^{\prime}_{T},c^{\prime}_{U})\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{m-t}$ and similarly consider $z^{\prime}=(x,y)\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{n-t}$. So it holds
@@
    For every possible pair of solutions $(\bar{x},\bar{y})$, compute $\bar{z}=A_{2}^{-1}(\bar{x},\bar{y})$.
In case in the decryption phase multiple solutions are obtained, we need to use special techniques (for example hash functions or redundancy in the plaintext) to make the decryption unique, see [DPS20a].
@@ 4.1.2 Proposal as a signature scheme
In a signature scheme, a sender produces a signature for a document, using the knowledge of the secret information. The receiver checks the validity of the signature for the received document.
@@
    We need to find $x\in{\mathbb{F}}_{q}^{t}$ and $y\in{\mathbb{F}}_{q}^{n-t}$ such that $G(x,y)=w^{\prime}$. So $T_{y}^{-1}(x)=w_{T}$ and $U(T_{y}^{-1}(x),y)=w_{U}$.
@@
Verification. The receiver wants to verify that the signature $(x^{\prime},y^{\prime})$ is valid for the document $d$.
@@
The correctness of the encryption scheme $G_{pub}(\bar{m})=c$ is verified in the same way.
@@
Regarding the function $U(x,y)$, different conditions must be satisfied. If we want to use $F$ as a central map for a signature scheme, we need that for any possible $(a,b)\in{\mathbb{F}}_{q}^{t}\times{\mathbb{F}}_{q}^{m-t}$, the system $U(a,y)=b$ always has a solution. Instead, for the encryption scheme, we need the system to have a few solutions. Indeed, in this case, the receiver has to consider all possible pre-images of the map $G_{pub}$ and find the correct one.
@@ 4.2 Our proposal
Assume that $F:{\mathbb{F}}_{q}^{n}\rightarrow{\mathbb{F}}_{q}^{m}$ is a quadratic function which admits a $t$-twist, so $F(x,y)=(T(x,y),U(x,y))$ where both $T$ and $U$ have at most degree 2. In what follows, we present our proposal for the choice of the maps $T$ and $U$.
@@ 4.2.1 The choice of $T$
leading to a $t$-twist of the form $G(x,y)=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. Observe that with this choice of $T$, to compute $\bar{x}=T_{\bar{y}}(w_{T})$ simply corresponds to computing $\bar{x}=w_{T}+{\mathfrak{q}}(\bar{y})$.
@@ 4.2.2 The choice of $U$
with $\{z_{j}:j\in V\}=\{x_{1},\ldots,x_{t},y_{1},\ldots,y_{s}\}$ and $\{z_{j}:j\in O\}=\{y_{s+1},\ldots,y_{n-t}\}$ respectively the sets of vinegar and oil variables, and with the coefficients $\alpha_{jk}^{(i)},\beta_{jk}^{(i)},\gamma_{j}^{(i)},\delta^{(i)}$ randomly chosen over ${\mathbb{F}}_{q}$. Notice that, fixed the vinegar variables, the system is linear in the oil variables, hence it is easy to solve, for example with a simple Gaussian reduction. The legitimate user can get the preimages of $U$ with respect to $y$ (fixed $x$) using classical techniques from OV systems [DY09, DPS20b]. Notice that the easiest system is obtained for $s=0$, however from the analysis presented in §6, this is not a good choice.
@@ 4.2.3 Pesto
Definition 4.2 (Pesto scheme).
@@
Set $G=(x-{\mathfrak{q}}(y),U(x-{\mathfrak{q}}(y),y))$. Then the map $G_{pub}=A_{1}\circ G\circ A_{2}$ is the public key, and $\langle A_{1},A_{2},{\mathfrak{q}},U\rangle$ constitutes the secret key.
In the following remark, we stress the role of $s$.
@@
Hence, when choosing the parameters $n,m,t,s$, one has also to consider if the linear system to solve, $m-t$ equations in $n-t-s$ variables, should be:
@@ 4.2.4 Toy Example
therefore $T(x,y)=x+{\mathfrak{q}}(y)$. The map $U$ is an OV system of 2 equations with $x_{1},x_{2},y_{1}$ vinegar variables and $y_{2},y_{3}$ oil variables. Hence we consider